OpenStudy (anonymous):

2x +5/10x + 3 = 1/3 solve for x

OpenStudy (anonymous):

2x +5/10x + 3 = 1/3 25/10x+3=1/3 25/10x=-8/3 x=-16/15

OpenStudy (anonymous):

where did 25 come from

OpenStudy (anonymous):

majic

OpenStudy (anonymous):

I made 2x into a fraction and found a common denominator to combine 2/1+5/10

OpenStudy (anonymous):

i dont know im sorry

OpenStudy (anonymous):

polpak could you help me with my question by chance?

OpenStudy (anonymous):

\[2x = \frac{2x*10x}{10x} = \frac{20x^2}{10x}\]

OpenStudy (anonymous):

let me take a moment to look it...take it in lol

OpenStudy (anonymous):

np, I was just correcting mm

OpenStudy (anonymous):

Just multiply both sides by 10x with the requirement that x cannot be 0

OpenStudy (anonymous):

polpak, do you mind showing me step by step

OpenStudy (anonymous):

Well I told you the first step.. what do you have when you multiply both sides by 10x?

OpenStudy (anonymous):

sorry...interrupted

OpenStudy (anonymous):

(2x+5)/3 = 10x/3

OpenStudy (anonymous):

where did the /3 on the left come from??

OpenStudy (anonymous):

wait.. is the original equation \[\frac{2x + 5}{10x + 3} = \frac{1}{3}\]

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

you need to use parens better when describing the problem ;p

OpenStudy (anonymous):

ok, so multiply both sides by 10x + 3 (with the stipulation that x cannot be -3/10)

OpenStudy (anonymous):

so far i have 2x + 4 = 10x/3...on the right track?

OpenStudy (anonymous):

no. you should have 2x + 5 = (10x + 3)/3

OpenStudy (anonymous):

i skipped to the next step...where 3/3 = 1 and then subtracted from other side...does x=1

OpenStudy (anonymous):

You can't do that

OpenStudy (anonymous):

Oh, I see

OpenStudy (anonymous):

Err wait. What did you do? x=1 isn't right

OpenStudy (anonymous):

4/5?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Show your work so I can figure out what you're doing

OpenStudy (anonymous):

so 2x + 5 = 10x +1 is that correct and i can go from there

OpenStudy (anonymous):

No

OpenStudy (anonymous):

You have to divide the whole 10x + 3 by 3, you can't just divide the 3 on top and leave the 10x alone. I suggest instead multiplying both sides by 3 to get rid of the fraction

OpenStudy (anonymous):

i know what i did...just noticed that...i started over with a clean piece of paper

OpenStudy (anonymous):

so 6x + 15 = 30x + 9

OpenStudy (anonymous):

no

OpenStudy (anonymous):

you forgot that the part on the right was divided by 3, so multiplying by 3 just cancels the denominator on the right side

OpenStudy (anonymous):

i left my brain behind...sorry

OpenStudy (anonymous):

np, it happens =)

OpenStudy (anonymous):

give me one minute..going to reason through...what bothers me the most...is that this looks simple and i feel like i can solve more difficult problems more easily

OpenStudy (anonymous):

\[3(2x + 5) = \frac{3(10x + 3)}{3}\]

OpenStudy (anonymous):

you should have that written on your paper

OpenStudy (anonymous):

so which 3's should cancel eachother out?

OpenStudy (anonymous):

\[\frac{3(10x + 3)}{3} = \frac{3}{3}(10x + 3) = 1(10x + 3) = 10x + 3\]

OpenStudy (anonymous):

ok...that was my problem....i wanted to cancel out the wrong 3...noted

OpenStudy (anonymous):

x = 3

OpenStudy (anonymous):

indeed

OpenStudy (anonymous):

finally lol....sorry this has been a chore...thanks for your help and patience

OpenStudy (anonymous):

np, just be sure to go back over the process and make sure you understand not only what we did, but why what you were doing doesn't work

OpenStudy (anonymous):

definitely will...starred it on my sheet as "needs more attention" lol

OpenStudy (anonymous):

good job =)

OpenStudy (anonymous):

thanks :)